# Federated Continual Learning Goes Online: Leveraging Uncertainty for Modality-Agnostic Class-Incremental Learning

### v2

@@ Abstract
Given the ability to model more realistic and dynamic problems, Federated Continual Learning (FCL) has been increasingly investigated recently. A well-known problem encountered in this setting is the so-called catastrophic forgetting, for which the learning model is inclined to focus on more recent tasks while forgetting the previously learned knowledge. The majority of the current approaches in FCL propose generative-based solutions to solve said problem. However, this setting requires multiple training epochs over the data, implying an offline setting where datasets are stored locally and remain unchanged over time. Furthermore, the proposed solutions are tailored for vision tasks solely. To overcome these limitations, we propose a new modality-agnostic approach to deal with the online scenario where new data arrive in streams of mini-batches that can only be processed once. To solve catastrophic forgetting, we propose an uncertainty-aware memory-based approach. In particular, we suggest using an estimator based on the Bregman Information (BI) to compute the model’s variance at the sample level. Through measures of predictive uncertainty, we retrieve samples with specific characteristics, and – by retraining the model on such samples – we demonstrate the potential of this approach to reduce the forgetting effect in realistic settings.
@@ 1 Introduction
In recent years, federated learning (FL) (McMahan et al., 2017) has received increasing attention for its ability to allow local clients work towards the same objective without the need for sharing limited and sensitive information. However, despite the benefits provided by the privacy-preserving collaborative training, the assumptions of the standard scenario are far from realistic (Babakniya et al., 2024). For example, the standard static-single-task framework has very limited practical applicability in real-world cases where local clients often need to continuously learn new tasks. Let us consider the problem of classifying new COVID-19 variants as illustrative use-case. Given the evolving nature of the virus, new variants (i.e., new classes) arise over time. In this context, healthcare facilities which collaborate using FL would fail since the standard setting does not consider the dynamic increment of new classes. For this reason, a new paradigm was recently introduced to model more complex dynamics: federated continual learning (FCL) (Yoon et al., 2021a; Dong et al., 2022). This new paradigm takes the global-local communication and privacy-preserving abilities enabled by FL and combines them with the Continual Learning (CL) ability of learning different tasks sequentially over time. In this way, FCL frameworks allow local clients to learn continuously from a stream of data where new classes can be added at any time. Consequently, this new paradigm not only shares the characteristics at the intersection of FL and CL, but also inherits their respective challenges. In particular one of the most prominent problems in CL, the so-called catastrophic forgetting for which the model is prone to suffer from significant performance degradation on older tasks. In CL, this problem is addressed in different ways ranging from memory-based approaches (Yoon et al., 2021b; Kumari et al., 2022; Hurtado et al., 2023) to generative methods (Shin et al., 2017; Lesort et al., 2019) or to regularization techniques (Lee et al., 2017; Aljundi et al., 2018; Chaudhry et al., 2018b; He and Jaeger, 2018). In the context of FCL, instead, the majority of the proposed approaches exploit generative models at both local and global level to trace and encode the past information and generate synthetic instances which faithfully mimic previous history whilst preserving data privacy. A substantial drawback of such generative approaches is that they are tailored to image data and it is not clear how they could translate to other data modalities.
@@
In realistic conditions however, all data from the current task may not be available for collection at the same time but may rather arrive in small chunks sequentially. Furthermore, given the ubiquity of smart and edge devices with limited capabilities (e.g., wearable devices), there is a need for updating with high frequency the learning model with the new incoming data to minimize memory overload and communication bandwidth (Ma et al., 2022). This problem of learning from an online stream of data is well investigated in CL while it remains largely unexplored in FCL.
@@
Inspired by the problem of training models over online streams of data, we first formalise a new scenario to tackle the online problem in FCL (online-FCL). In this new scenario, we assume each client to learn from a stream of data where new data arrive in mini-batches which can only be processed once. As such, in line with the definition of online-CL, the model is updated with high frequency (Soutif-Cormerais et al., 2023). Then, taking inspiration from the most popular solutions in online-CL, we introduce memory buffers to alleviate catastrophic forgetting at local level. More in detail, we propose an uncertainty-based memory management where data points are stored in local buffers according to their predictive uncertainty. Intuitively, predictive uncertainty provides a glimpse of the samples’ location in the decision space; samples with low uncertainty are the most representative ones for the respective class, while samples with high uncertainty represent data points that are close to the decision boundary and/or outliers. Thus, via estimates of predictive uncertainty, we can store in the memory samples with desired properties. Here, we propose to quantify uncertainty by directly estimating a generalized variance term from the loss function, which can be interpreted as a measure of epistemic uncertainty. To this end, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss (Gruber and Buettner, 2023) and estimate the Bregman Information (BI) as the variance term in logit space. The uncertainty-based memory management makes our approach modality-agnostic. In fact, independent of the data modality (e.g., images, texts), our solution allows us to compute uncertainty estimates and thus populate the memory accordingly. In contrast to most of the current solutions in FCL which are limited to vision tasks and the offline setting, our simple but effective baseline is flexible in terms of data modality.
@@
In the last part of the paper, we demonstrate the ability of the proposed approach to reduce the forgetting effect in different scenarios. In the first part of the experiments, we evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009), a standard dataset used in this context. The goal is to understand how the proposed (epistemic) uncertainty estimate based on the Bregman Information performs in comparison with other standard estimates of overall model uncertainty under a memory-based regime. Then, departing from the common evaluation pipeline that would involve datasets like EMNIST (as in Qi et al. (2023); Wuerkaixi et al. (2024)) or larger datasets from the same naturalistic domain (e.g., CIFAR100 or ImageNet – as in Qi et al. (2023); Babakniya et al. (2024)), we validate our results on more probing real-world datasets from the medical domain. Finally, to showcase the ability of our approach to work with different data modalities, we test our findings on a text classification task.
@@
  • We propose and formalise a novel framework to tackle the FCL problem in the online setting.
  • We highlight the limitations of the current state-of-the-art generative-based solution to work in the online setting.
  • We propose a modality-agnostic memory-based solution that employs an alternative estimate for predictive uncertainty, which stems directly from a bias-variance decomposition of the cross-entropy loss for classification tasks (Gruber and Buettner, 2023), to populate the memory.
  • We demonstrate the efficacy of our method in more realistic scenarios including datasets from different domains, imbalanced, and with different modalities.
Figure 1: Schematic overview of the proposed online-FCL scenario.
@@ 2 Related Work
Here we briefly summarise the identified limitations of such generator-based approaches. As anticipated before, they assume to work in the offline setting where static datasets are collected in advance and stored locally. This allows them to be trained over the same task dataset many times which is needed to reach convergence and to learn meaningful patterns in order to generate meaningful synthetic images. Finally, the storage of the whole task datasets and the generator models may be unfeasible in resource-limited devices.
3 Our Approach: O-FCIL
@@ 3.2 Methodology
In the following subsections, we describe in more detail the characteristics of our approach at the client and server levels. For the client side, we outline our uncertainty-aware memory management detailing the properties of the employed uncertainty estimate and its advantages compared to standard scores. For the server side, we detail the adjustments done to improve the communication and parameter averaging effectiveness in the online scenario.
@@ 3.2.1 Client-level
Different measure of predictive uncertainty can capture distinct aspects of a model’s irreducible aleatoric uncertainty (inherent in the data) and its epistemic uncertainty (that stems from the finite training data and can be reduced by gathering more data). In online CL, the most commonly used measures are derived directly from the confidence scores of the model and mostly capture the irreducible aleatoric uncertainty (Wimmer et al., 2023). Here, we hypothesize it may be more beneficial for the model to replay instances which are representative in the sense that there is low uncertainty about the data generating process (we refer to this as low epistemic uncertainty, while acknowledging that there exist varying definitions). To not rely on specialized Bayesian models, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss and compute a generalized variance term directly from the loss (Gruber and Buettner, 2023). Such bias-variance decomposition decomposes the expected prediction error (loss) into the model’s bias, variance, and an irreducible error (noise term). The latter is related to aleatoric uncertainty, whereas the variance term can directly be related to epistemic uncertainty (Gruber et al., 2023).
@@ 3.2.2 Global-level
Although clients work on the same task, they may receive data coming from different classes (as depicted in Figure 1). However, it may happen that some of them receive mini-batches that contain the same label (see $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ in Figure 1). Thus, we need to consider that local parameters collected for parameter averaging can be biased towards the most common classes. For this, we propose to first create an aggregated model for each class available in the current round, and then compute $\bm{\theta}_{g}^{r}$ using the class-based models just created. In this way, the current classes contribute equally during the updates of the global parameters. In case all the clients work on the same task sequence, the parameter averaging results in the standard computation, like e.g., FedAvg (McMahan et al., 2017).
Finally, once the new global parameters $\bm{\theta}_{g}^{r}$ are computed, to avoid a drastic change between the old parameters and the new ones, following Shenaj et al. (2023), we propose to average the newly computed parameters at round $r$ with the previous parameters computed at round $r-1$.
@@ 4.1 Datasets and Settings
To understand the behaviour of different algorithms in the online-FCL scenario, we use a dataset commonly used in the literature, namely, CIFAR10 (Krizhevsky et al., 2009). We randomly assign a set of classes to 5 tasks and we split the task data evenly among the clients such that the task sequence $T_{k}^{t}$ for task $t$ and client $k$ has two classes and does not share any data with the other clients. In this way, since the class assignment per task is different every time, we can identify the strategy providing greater flexibility and effectiveness irrespective of the composition of the tasks. Then, as anticipated in the introductory part, we assess the performance under more difficult and realistic conditions. Instead of using datasets that, to some extent, share similar characteristics with CIFAR10 (i.e., CIFAR100 and ImageNet) or represent unrealistic tasks (such as EMNIST), we validate our results on datasets for biomedical image analysis. Apart from the change in the domain (which in turn means different backgrounds in the images, different statistics, etc.), these datasets pose an additional challenge compared to the standard benchmark datasets, i.e., data imbalance. To reflect realistic conditions where recent tasks contain fewer data points than the older ones since the time to collect them is shorter, we assign classes to tasks based on the class size. We decide to focus on two biomedical datasets annotated by expert clinical pathologists; the colorectal cancer hystology (CRC-Tissue) dataset (Kather et al., 2019; Yang et al., 2023) containing images divided in 8 classes of hematoxylin–eosin (HE)–stained slides taken from patients with colorectal cancer (CRC), and the kidney cortex cells (KC-Cell) dataset containing 8 classes of human kidney cortex tissue sections (Ljosa et al., 2012; Yang et al., 2023). Finally, to test our approach on text classification tasks, we use the 20NewsGroups dataset (Lang, 1995) consisting of thousand of news evenly partitioned across 20 different newsgroups. As for CIFAR10, we randomly assign classes to tasks every run. Statistics and more details about the used datasets can be found in the supplementary material (Appendix A.3).
@@
Given the novelty of the online setting, there are no direct competitors which we could refer to. We therefore investigate the most successful algorithms from the online-CL and the FCL literature. We compare our approach with a standard baseline for FL, i.e., FedAvg (McMahan et al., 2017), a standard memory-based approach for online-CL, namely Experience Replay (ER) (Chaudhry et al., 2019), and the state-of-the-art approach for FCL, i.e., MFCL (Babakniya et al., 2024). Our decision to use ER is based on the fact that, despite its simplicity, it is surprisingly competitive when compared with more sophisticated and newer approaches as shown in recently conducted empirical surveys (Soutif-Cormerais et al., 2023). For FCL, we use MFCL because its data-free solution trains the generator on the server side. As such, we believe it may be potentially feasible in the online scenario, in contrast to other approaches that train local generators at the client-side (Qi et al., 2023). In addition to the standard ER, we also include the class-balanced version (CBR) presented in Chrysakis and Moens (2020). Finally, to show the competitiveness of $u(x)_{BI}$, we also consider other uncertainty scores. In particular, least confidence (LC), margin sampling (MS), ratio of confidence (RC), and entropy (EN) (Shannon, 1948; Campbell et al., 2000; Culotta and McCallum, 2005).
@@
In all the experiments for image classification, we employ a slim version of Resnet18 (He et al., 2016) – as done in previous works in online-CL (Kumari et al., 2022; Hurtado et al., 2023; Soutif-Cormerais et al., 2023) –, and use the SGD optimizer with a learning rate of $0.1$. For text classification, we employ a simple MLP with a single fully connected hidden layer (512 units) and Adam optimizer with a learning rate of $0.01$. To generate the perturbed inputs for estimating predictive uncertainty, we use two different strategies according to the different data modalities. For vision tasks, the perturbations are generated via standard augmentation. The list of augmentations used in our experiments is provided in the appendix. For the natural language experiment, we leverage recent progress in foundation models and first create general-purpose latent representations of the input texts via a pre-trained sentence embedder. In particular, we employ e5-small-v2 (Wang et al., 2022) (384 dimensions) via HuggingFace (Wolf et al., 2020). Then, for perturbing the generated vector representations, we add Gaussian noise from $\mathcal{N}(0,0.1)$ to each latent dimension.
@@
Following standard practice, we set the batch size equal to $10$. The memory size was set to different values for each dataset in order to examine the performance of various memory configurations, including both large and small buffers. The burn-in period is set to $30$. A communication round with the server is performed after $q=5$ mini-batches. The number of clients is set to $5$. For parameter averaging, we employ FedAvg (McMahan et al., 2017). Given the flexibility of the approach, the model averaging strategy can be changed as desired. For evaluation, following recent works (Yoon et al., 2021a; Wuerkaixi et al., 2024), we use the average last accuracy (A) and average forgetting (F) (see appendix A.4 for the definitions). All the experiments are run on three different random seeds. For each dataset, experiments were run on a Linux machine using a single Quadro RTX 5000 and 16 GB RAM.
@@
Figure 3: (left) Average accuracy per task at the end of the training in comparison with FL (FedAvg) and generative-based FCL baselines (MFCL); (right) Relative CF improvement over the CL baseline (ER).
4.2 Empirical Results
From the experiments conducted on CIFAR10, we can observe that the proposed approach is able to reduce CF consistently across different memory sizes and class-per-task assignments. The results in Table 1 suggest that storing the class-representative samples (i.e., the least uncertain data points for each class) provides a benefit in terms of predictive performance gain and forgetting reduction. Our findings are substantiated in the classification tasks for biomedical images. Although the images originate from another domain, Tables 2 and 3 confirm that in terms of CF, BI outperforms the considered baselines in all cases. Crucially, we also show the ability of our approach to work with imbalanced real-world data. Figure 3 summarise our findings; in comparison with standard baselines, our simple approach is able to perform consistently across tasks (left plot) and reduce CF on different datasets (right plot). Finally, we demonstrate that our method can work on other tasks than image classification. Table 4 reports the results on the 20NewsGroup dataset.
More generally, we can notice that the difference in terms of predictive accuracy is less marked. This may be due to the fact that the baselines have higher accuracy on the last tasks and poor performance on the first ones, while in our case the accuracy is kept more uniform across the tasks (reflected in a lower average forgetting) as shown in the left plot in Figure 3.
@@ 4.2 Empirical Results
Table 2: Comparison of average last accuracy (A) and last forgetting (F) on CRC-Tissue (4 tasks).

### v3

# Federated Continual Learning Goes Online: Uncertainty-Aware Memory Management for Vision Tasks and Beyond

@@ Abstract
Given the ability to model more realistic and dynamic problems, Federated Continual Learning (FCL) has been increasingly investigated recently. A well-known problem encountered in this setting is the so-called catastrophic forgetting, for which the learning model is inclined to focus on more recent tasks while forgetting the previously learned knowledge. The majority of the current approaches in FCL propose generative-based solutions to solve said problem. However, this setting requires multiple training epochs over the data, implying an offline setting where datasets are stored locally and remain unchanged over time. Furthermore, the proposed solutions are tailored for vision tasks solely. To overcome these limitations, we propose a new approach to deal with different modalities in the online scenario where new data arrive in streams of mini-batches that can only be processed once. To solve catastrophic forgetting, we propose an uncertainty-aware memory-based approach. Specifically, we suggest using an estimator based on the Bregman Information (BI) to compute the model’s variance at the sample level. Through measures of predictive uncertainty, we retrieve samples with specific characteristics, and – by retraining the model on such samples – we demonstrate the potential of this approach to reduce the forgetting effect in realistic settings while maintaining data confidentiality and competitive communication efficiency compared to state-of-the-art approaches.
@@ 1 Introduction
In recent years, federated learning (FL) (McMahan et al., 2017) has received increasing attention for its ability to allow local clients work towards the same objective without the need for sharing limited and sensitive information. However, despite the benefits provided by the privacy-preserving collaborative training, the assumptions of the standard scenario are far from realistic (Babakniya et al., 2024). For example, the standard static-single-task framework has very limited practical applicability in real-world cases where local clients often need to continuously learn new tasks. Let us consider the problem of classifying new COVID-19 variants as illustrative use-case. Given the evolving nature of the virus, new variants (i.e., new classes) arise over time. In this context, healthcare facilities which collaborate using FL would fail since the standard setting does not consider the dynamic increment of new classes. For this reason, a new paradigm was recently introduced to model more complex dynamics: federated continual learning (FCL) (Yoon et al., 2021a; Dong et al., 2022). This new paradigm takes the global-local communication and privacy-preserving abilities enabled by FL and combines them with the Continual Learning (CL) ability of learning different tasks sequentially over time. However, FCL not only shares the characteristics at the intersection of FL and CL, but also inherits their respective challenges. In particular one of the most prominent problems in CL, the so-called catastrophic forgetting for which the model is prone to suffer from significant performance degradation on older tasks. In CL, this problem is addressed in different ways ranging from memory-based approaches (Yoon et al., 2021b; Kumari et al., 2022; Hurtado et al., 2023) to generative methods (Shin et al., 2017; Lesort et al., 2019) or to regularization techniques (Lee et al., 2017; Aljundi et al., 2018; Chaudhry et al., 2018b; He and Jaeger, 2018). In the context of FCL, instead, the majority of the proposed approaches exploit generative models at both local and global level to trace and encode the past information and generate synthetic instances which faithfully mimic previous history whilst preserving data privacy. A substantial drawback of such generative approaches is that they are tailored to image data and it is not clear how they could translate to other data modalities.
@@
In realistic conditions however, all data from the current task may not be available for collection at the same time but may rather arrive in small chunks sequentially. Furthermore, given the ubiquity of smart and edge devices with limited capabilities (e.g., wearable devices), there is a need for updating the learning model with the new incoming data to minimize memory overload and communication bandwidth (Ma et al., 2022). This problem of learning from an online stream of data is well investigated in CL while it remains largely unexplored in FCL.
@@
Inspired by the problem of training models over online streams of data, we first formalise a new scenario to tackle the online problem in FCL (online-FCL). In this new scenario, we assume each client to learn from a stream of data where new data arrive in mini-batches which can only be processed once. As such, in line with the definition of online-CL, the model is updated with high frequency (Soutif-Cormerais et al., 2023). Then, taking inspiration from the most popular solutions in online-CL, we introduce memory buffers to alleviate CF at local level. More in detail, we propose an uncertainty-based memory management where data points are stored in local buffers according to their predictive uncertainty. Intuitively, predictive uncertainty provides a glimpse of the samples’ location in the decision space; samples with low uncertainty are the most representative ones for the respective class, while samples with high uncertainty represent data points that are close to the decision boundary and/or outliers. Thus, via estimates of predictive uncertainty, we can store in the memory samples with desired properties. Here, we propose to quantify uncertainty by directly estimating a generalized variance term from the loss function, which can be interpreted as a measure of epistemic uncertainty. To this end, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss (Gruber and Buettner, 2023) and estimate the Bregman Information (BI) as the variance term in logit space.
The uncertainty-based memory management makes our approach flexible in terms of data modality. In fact, independent of the data modality (e.g., images, texts), our solution allows us to compute uncertainty estimates and thus populate the memory accordingly. This is in contrast to most of the current solutions in FCL which are limited to vision tasks and the offline setting.
@@
In the last part of the paper, we demonstrate the ability of the proposed approach to reduce the forgetting effect in different scenarios. In the first part of the experiments, we evaluate our approach on CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009), two standard datasets used in this context. The goal is to understand how the proposed (epistemic) uncertainty estimate based on the Bregman Information performs in comparison with other standard estimates of overall model uncertainty under a memory-based regime. Then, departing from the common evaluation pipeline that would involve datasets like EMNIST (as in Qi et al. (2023); Wuerkaixi et al. (2024)) or larger datasets from the same naturalistic domain (e.g., ImageNet – as in Qi et al. (2023); Babakniya et al. (2024)), we validate our results on more probing real-world datasets from the medical domain. Finally, to showcase the ability of our approach to work with different data modalities, we test our findings on text classification tasks.
@@
  • We propose and formalise a novel framework to tackle the FCL problem in the online setting which is largely overlooked in the literature.
  • We highlight the limitations of current state-of-the-art generative-based solutions to work in the online setting and show empirically their inefficiency in the experimental section.
  • We propose a memory-based solution that employs an alternative estimate for predictive uncertainty – which stems directly from a bias-variance decomposition of the cross-entropy loss for classification tasks – to populate the memory.
  • We demonstrate the efficacy of our method in more realistic scenarios including datasets from different domains, with imbalance, and with different modalities.
Figure 1: Schematic overview of the proposed online-FCL scenario. Each client $\mathcal{C}_{k}$ receives their data as a stream of batches. Each batch $b_{k}^{t}$ is used for training and for updating the local memory $\mathcal{M}_{k}$. Once the current batch is processed, the client receives a new batch and continues their training without the possibility of retraining on previously seen batches.
@@ 2 Related Work
Here we briefly summarise the identified limitations of such generator-based approaches. As anticipated before, they assume to work in the offline setting where static datasets are collected in advance and stored locally. This allows them to be trained over the same task dataset many times which is needed to reach convergence and to learn meaningful patterns in order to generate informative synthetic images. Finally, the storage of the whole task datasets and the generator models may be unfeasible in resource-limited devices.
3 Our Approach: O-FCL
@@ 3.2 Methodology
In the following subsections, we describe in more detail the characteristics of our approach at the client and server levels. For the client side, we outline our uncertainty-aware memory management detailing the properties of the employed uncertainty estimate and its advantages compared to standard scores. For the server side, we detail the adjustments done to improve the communication and parameter averaging effectiveness in the online scenario. A summary of the complete training procedure is reported in Algorithm 1 (see Appendix A.9).
@@ 3.2.1 Client-level
Different measure of predictive uncertainty can capture distinct aspects of a model’s irreducible aleatoric uncertainty (inherent in the data) and its epistemic uncertainty (that stems from the finite training data and can be reduced by gathering more data). In online CL, the most commonly used measures are derived directly from the confidence scores of the model and mostly capture the irreducible aleatoric uncertainty (Wimmer et al., 2023). Here, we hypothesize it may be more beneficial for the model to replay instances which are representative in the sense that there is low uncertainty about the data generating process (we refer to this as low epistemic uncertainty, while acknowledging that there exist varying definitions). To not rely on specialized Bayesian models, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss and compute a generalized variance term directly from the loss (Gruber and Buettner, 2023). Such bias-variance decomposition decomposes the expected prediction error (loss) into the model’s bias, variance, and an irreducible error (noise term). The latter is related to aleatoric uncertainty, whereas the variance term can directly be related to epistemic uncertainty (Gruber et al., 2023). This is the first attempt of using this decomposition for memory management in FCL contexts.
@@ 3.2.2 Global-level
Although clients work on the same task, they may receive data coming from different classes where some of them are over-represented compared to others. In this case, the local parameters collected for parameter averaging can be biased towards the most common classes. For this, we suggest to first create an aggregated model for each class available in the current round, and then compute $\bm{\theta}_{g}^{r}$ using the class-based models just created. This way, the current classes contribute equally during the updates of the global parameters. In case all the clients receive data containing classes equally distributed, the parameter averaging results in the standard computation, like e.g., FedAvg (McMahan et al., 2017). Importantly, if sharing the class information is not possible, we can rely on standard averaging strategies (e.g., FedAvg or FedProx (Li et al., 2020)) without hampering the performance since our methodology is flexible on this matter (see Table 8). Finally, once the new global parameters $\bm{\theta}_{g}^{r}$ are computed, to avoid a drastic change between the old parameters and the new ones, following Shenaj et al. (2023), we propose to average the newly computed parameters at round $r$ with the previous parameters computed at round $r-1$.
@@ 4.1 Datasets and Settings
To understand the behaviour of different algorithms in the online-FCL scenario, we use two datasets commonly used in the literature, namely, CIFAR10 and CIFAR100 (Krizhevsky et al., 2009). We randomly assign a set of classes to 5 tasks and we split the task data evenly among the clients such that the task sequence $T_{k}^{t}$ for task $t$ and client $k$ does not share any data with the other clients. In this way, since the class assignment per task is different every time, we can identify the strategy providing greater flexibility and effectiveness irrespective of the composition of the tasks. Then, as anticipated in the introductory part, we assess the performance under more difficult and realistic conditions. Instead of using datasets that, to some extent, share similar characteristics with CIFAR (i.e., ImageNet) or represent unrealistic tasks (such as EMNIST), we validate our results on datasets for biomedical image analysis. Apart from the change in the domain (which in turn means different backgrounds in the images, different statistics, etc.), these datasets pose an additional challenge compared to the standard benchmark datasets, i.e., data imbalance. To reflect realistic conditions where recent tasks contain fewer data points than the older ones since the time to collect them is shorter, we assign classes to tasks based on the class size. We decide to focus on two biomedical datasets annotated by expert clinical pathologists; the colorectal cancer histology (CRC-Tissue) dataset (Kather et al., 2019; Yang et al., 2023) containing images divided in 8 classes of hematoxylin–eosin (HE)–stained slides taken from patients with colorectal cancer (CRC), and the kidney cortex cells (KC-Cell) dataset containing 8 classes of human kidney cortex tissue sections (Ljosa et al., 2012; Yang et al., 2023). Finally, to evaluate our approach on text classification tasks, we employ three different datasets: the 20NewsGroups dataset (Lang, 1995), DBPedia (Zhang et al., 2015), and Yahoo Answers (Zhang et al., 2015). As for vision tasks, we randomly assign classes to tasks every run. Statistics and more details about the used datasets can be found in the supplementary material (Appendix A.3).
@@
Given the novelty of the online setting, there are no direct competitors which we could refer to. We therefore investigate the most successful algorithms from the online-CL and the FCL literature. We compare our approach with standard baselines for FL, i.e., FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020), a standard memory-based approach for online-CL, namely Experience Replay (ER) (Chaudhry et al., 2019), and two state-of-the-art approaches for FCL with generative replay, namely MFCL (Babakniya et al., 2024) and FedCIL (Qi et al., 2023). Our decision to use ER is based on the fact that, despite its simplicity, it is surprisingly competitive when compared with more sophisticated and newer approaches as shown in recently conducted empirical surveys (Soutif-Cormerais et al., 2023). For FCL, we are particularly interested in MFCL because its data-free solution trains the generator on the server side. As such, we believe it may be potentially feasible in the online scenario, in contrast to other approaches such as FedCIL that train local generators at the client-side (Qi et al., 2023). In addition to the standard ER, we also include the class-balanced version (CBR) presented in Chrysakis and Moens (2020). Finally, to show the competitiveness of $u(x)_{BI}$, we also consider other uncertainty scores. In particular, least confidence (LC), margin sampling (MS), ratio of confidence (RC), and entropy (EN) (Shannon, 1948; Campbell et al., 2000; Culotta and McCallum, 2005).
@@
In all the experiments for image classification, we employ a slim version of Resnet18 (He et al., 2016) – as done in previous works in online-CL (Kumari et al., 2022; Hurtado et al., 2023; Soutif-Cormerais et al., 2023) –, and use the SGD optimizer with a learning rate of $0.1$. For text classification, we employ a simple MLP with a single fully connected hidden layer (512 units) and Adam optimizer with a learning rate of $0.01$. To generate the perturbed inputs for estimating predictive uncertainty, we use two different strategies according to the different data modalities. For vision tasks, the perturbations are generated via standard augmentation. The list of augmentations used in our experiments is provided in Appendix A.2. For the natural language experiments, we leverage recent progress in foundation models and first create general-purpose latent representations of the input texts via a pre-trained sentence embedder. In particular, we employ e5-small-v2 (Wang et al., 2022) (384 dimensions) via HuggingFace (Wolf et al., 2020). Then, for perturbing the generated vector representations, we add Gaussian noise from $\mathcal{N}(0,0.1)$ to each latent dimension.
@@
Following standard practice, we set the batch size equal to $10$. The memory size was set to different values for each dataset in order to examine the performance of various memory configurations, including both large and small buffers. The burn-in period is set to $30$. A communication round with the server is performed after $q=5$ mini-batches. The number of clients is set to $5$. For parameter averaging, we employ FedAvg (McMahan et al., 2017). Given the flexibility of the approach, as mentioned before, the model averaging strategy can be changed as desired. For evaluation, following recent works (Yoon et al., 2021a; Wuerkaixi et al., 2024), we use the average last accuracy (A) and average forgetting (F) (see appendix A.4 for the definitions). All the experiments are run on three different random seeds. For each dataset, experiments were run on a Linux machine using a single Quadro RTX 5000 and 16 GB RAM.
@@
Figure 3: (left) Average accuracy per task in comparison with FL (FedAvg) and generative-based FCL baselines (MFCL); (right) Relative CF improvement over the CL baseline (ER).
@@
4.2 Empirical Results
From the experiments conducted on the CIFAR datasets, we can observe that the proposed approach is able to reduce CF consistently across different memory sizes and class-per-task assignments. The results in Tables 1 and 2 suggest that storing the class-representative samples (i.e., the least uncertain data points for each class) provides a benefit in terms of predictive performance gain and forgetting reduction. Our findings are substantiated in the classification tasks for biomedical images. Although the images originate from another domain, Table 3 confirms that in terms of CF, BI outperforms the considered memory-based baselines in most of the cases. Crucially, we also show the ability of our approach to work with imbalanced real-world data. Table 5 and Figure 3 summarise our findings; in comparison with both memory-based and generative-based methods, our simple-yet-elegant approach is able to perform consistently across tasks (left plot) and reduce CF on different datasets (right plot). Finally, we demonstrate that our method outperforms baselines also on non-vision data. Table 4 reports the results on the textual classification datasets. Here we find that the difference in terms of predictive accuracy is less marked. This is because the baselines have higher accuracy on the last tasks and poor performance on the first ones, while in our case the accuracy is kept more uniform across the tasks (reflected in a lower average forgetting) as shown in the left plot in Figure 3. Note that the use of generative baselines is restricted to image data only.
Table 2: Comparison of last accuracy (A) and last forgetting (F) on CIFAR100 (5 tasks).
Score M=1000 M=2000
ER 10.33 $\pm$ 0.58 10.95 $\pm$ 0.4
6.89 $\pm$ 1.53 9.12 $\pm$ 0.96
CBR 13.38 $\pm$ 0.58 11.41 $\pm$ 0.36
31.66 $\pm$ 0.30 9.37 $\pm$ 1.56
Top Bottom Top Bottom
LC A()($\uparrow$)( ↑ ) 12.71 $\pm$ 0.96 13.43 $\pm$ 1.00 13.50 $\pm$ 0.81 14.09 $\pm$ 0.47
F()($\downarrow$)( ↓ ) 8.03 $\pm$ 1.88 6.95 $\pm$ 0.91 7.99 $\pm$ 1.34 7.75 $\pm$ 1.87
MS A()($\uparrow$)( ↑ ) 13.34 $\pm$ 1.71 13.77 $\pm$ 1.31 13.83 $\pm$ 0.56 14.26 $\pm$ 0.72
F()($\downarrow$)( ↓ ) 7.48 $\pm$ 2.04 6.84 $\pm$ 1.60 8.04 $\pm$ 0.99 8.01 $\pm$ 1.76
RC A()($\uparrow$)( ↑ ) 13.30 $\pm$ 1.47 13.76 $\pm$ 1.24 13.28 $\pm$ 1.23 13.83 $\pm$ 0.69
F()($\downarrow$)( ↓ ) 8.20 $\pm$ 1.91 6.97 $\pm$ 1.12 7.52 $\pm$ 1.70 7.70 $\pm$ 1.58
EN A()($\uparrow$)( ↑ ) 12.77 $\pm$ 1.00 12.77 $\pm$ 0.71 13.81 $\pm$ 1.00 13.93 $\pm$ 0.76
F()($\downarrow$)( ↓ ) 8.11 $\pm$ 1.14 7.90 $\pm$ 1.20 8.10 $\pm$ 1.51 7.50 $\pm$ 1.17
BI A()($\uparrow$)( ↑ ) 12.89 $\pm$ 0.97 14.04 $\pm$ 0.62 13.72 $\pm$ 0.70 14.31 $\pm$ 0.76
F()($\downarrow$)( ↓ ) 7.54 $\pm$ 1.43 6.73 $\pm$ 1.76 8.06 $\pm$ 1.49 6.96 $\pm$ 1.40
Table 3: Comparison of average last accuracy (A) and last forgetting (F) on (left) CRC-Tissue (4 tasks), and (right) KC-Cell (4 tasks).
CRC-Tissue